Solve for the zeroes: x^2+5x-6
Answer by jim_thompson5910(35256)

x^2+5x-6 = 0

(x+6)(x-1) = 0

x+6=0 or x-1 = 0

x = -6 or x = 1

So the two solutions are x = -6 or x = 1